1 ounce of the mixture containing 6% salt is to be mixed with 2 ounces of a mixture which is 15% salt, in order to obtain a solution which is 12% salt. What is the percentage of salt in the resulting solution?
when i try to do it on this website, it gives a syntax error on the last step so i can't see the answer spmebody please tell me the answer please.

Answer by stanbon(75887)

1 ounce of the mixture containing 6% salt is to be mixed with 2 ounces of a mixture which is 15% salt, in order to obtain a solution which is 12% salt. What is the percentage of salt in the resulting solution?
----------------------
6% solution DATA;
Amt = 1 oz ;Amt of salt = 0.06 oz
-------------------------
15% solution DATA;
Amt = 2 oz ; Amt of salt = 0.15*2 = 0.30 oz
----------------------------
12% mixture DATA:
Amt = 3 oz ; Amt of salt = 0.12*3 = 0.36 oz
-----------------------------
EQUATION:
Amt of mixtue: 3 oz
Amt of salt in the mixture: 0.36 oz
Percent of salt in mixture = 0.36/3 = 0.12 = 12%
=================================
Cheers,
Stan H.